OpenStudy (anonymous):

What is the point-slope equation of line with slope -3 that contains the point (-8,-4)?

OpenStudy (chosenmatt):

WHALECOME TO OPENSTUDY

OpenStudy (anonymous):

Point slope form \[y-y_1 = m(x-x_1)\] (x1,y1) = given point m = slope

OpenStudy (chosenmatt):

lol

OpenStudy (anonymous):

thank you.